chris lynch plans to invest $200 into a money market account. Find the interest rate that is needed for the money to grow $1,400 in 18 years if the interest is compounded quarterly.
Answer by Boreal(15235)

P=Po{1+r/n)^nt
1400=200{1+r/4)^72
7={1+r/4)^72
ln of both sides
1.9459=72 ln {1+r/4}
0.02703=ln (1+r/4)
raise to e power
1.0274=1+r/4
.0274=r/4
r=10.96%.
Check with 11% and see if Po increases 7 fold
{1+11/4)}^72
1.0275^72=7.0516 This is a little more than 7, but I used 11% and not 10.96%